THE ROVIBRATIONAL SPECTRUM OF THE RING PUCKERING AND NH BENDING MODES OF AZE-TIDINE

T. ZAPOROZAN and \underline{J. A. VAN WIJNGAARDEN}, Department of Chemistry, University of Manitoba, Winnipeg, MB, Canada R3T 2N2.

High quality infared spectra of the two lowest frequency modes of azetidine (c-C_{3}H_{6}NH) have been recorded between 100 and 700 cm^{−1} using the far infrared beamline at the Candian Light Source in Saskatoon. The far infrared endstation couples highly spatially confined synchrotron light into a Bruker IFS125HR Fourier transform spectrometer and can achieve spectral resolution up to 0.000959 cm^{−1}. The observed modes correspond to the ring puckering mode (ν_{16}) at 207.2 cm^{−1} and the N-H bending mode (ν_{15}) at 648.2 cm^{−1}. Rotational analysis of the ring puckering mode confirmed our ab initio prediction that the band structure is governed by a−type selection rules rather than c−type transitions as reported for oxetane.^{a} The results of the rovibrational analysis of both bands will be presented and comparisons will be made between the synchrotron spectra and those collected using a conventional globar source.

aM. Winnewisser, M. Kunzmann, M. Lock, and B. P. Winnewisser, J. Mol. Struct. 561, 1 (2001).